Complete the following problems and round the answers to the correct number of significant figures.
(i). (23.4g + 102.4g + 0.003g)/6.478mL (ii). (74.335m – 73.332m)/(4.75s X 1.114s)
1
Expert's answer
2020-05-03T15:35:23-0400

(i). (23.4g + 102.4g + 0.003g)/6.478mL

23.4 + 102.4 = 125.8 (g)

125.8 + 0.003 = 125.8 (g)

125.8 / 6.478 = 19.42 or 1.942 × 101(g/mL)


 (ii). (74.335m – 73.332m)/(4.75s X 1.114s)

74.335 - 73.332 = 1.003 (m)

4.75 * 1.114 = 5.29 (s)

1.003 / 5.29 = 0.190 or 1.90 × 10-1(m/s)